determine the number of atoms in each of the following
1) 2.85 mol of Ag
2) 12.9g of CaCO3
3) 12.9g of Fe(NO3)3
1
Expert's answer
2012-08-09T08:27:42-0400
The number of particles is
N(particles) =
N(Avogadro)*n(particles)
1)
N(Ag) = 6.022*10^23 mol^-1 * 2.85 mol =
1.72*10^24
2)
The amount of CaCO3 is
n(CaCO3) = m(CaCO3) / M(CaCO3) =
12.9 g / 100 g/mol = 0.129 mol
N(CaCO3) = 6.022*10^23 mol^-1 * 0.129 mol =
7.77*10^22
3)
The amount of Fe(NO3)3 is
n(Fe(NO3)3) = m(Fe(NO3)3) /
M(Fe(NO3)3) = 12.9 g / 242 g/mol = 0.053 mol
N(Fe(NO3)3) = 6.022*10^23 mol^-1
* 0.053 mol = 3.19*10^22